# Beyond Pipelines: A Survey of the Paradigm Shift toward Model-Native Agentic AI

### v1

@@ 5.1. Overview
To systematically review this evolution, this section divides memory into two primary layers: short-term memory and long-term memory. Our discussion of short-term memory follows its own developmental trajectory. We begin with techniques for long context, which address the fundamental challenge of processing large volumes of information within a single session (Beltagy et al., 2020; Dao et al., 2022; Zaheer et al., 2020). We then examine context management, a more advanced capability focused on the selective organization and utilization of information to mitigate attention dilution and improve reasoning quality (Asai et al., 2023). In contrast, the section on long-term memory reviews methods for persistently storing knowledge across sessions, either in external repository or within model parameters (Wang et al., 2024c; Xu et al., 2025a; Chandy, 2025; Zhong et al., 2023). Notably, since long-term memory content is typically not rewritable online, the paradigm shift in this area focuses more on internalizing retrieval and usage strategies rather than the entire storage process.
@@ 5.2.1. Pipeline-based Paradigm
When the dialogue context far exceeds the model’s budget and the necessary knowledge resides in external stores, RAG dynamically constructs the working context for each turn. Compared with traditional RAG pipelines, which typically treat retrieval as a stateless, per-query operation, conversational variants are session-aware: they employ query expansion to capture dialogue intent and multi-path retrieval with reranking to optimize the evidence set across turns. This approach excels at knowledge-intensive, needle-in-a-haystack problems like log analysis and technical question answering. Various frameworks have demonstrated this approach, from those employing MapReduce paradigms (Harrison, 2 10) to more advanced systems that utilize graph-based structures (Edge et al., 2025) or memory-augmented techniques (Qian et al., 2025b; Tang et al., 2024). A key challenge remains that excessive injection of retrieved information can lead to attention dilution. Therefore, practical deployments often treat long windows and session RAG as complementary, using the window to cover retrieval misses and falling back to precise injection when the window becomes too noisy.
@@ 5.2.2. Model-native Paradigm
Expanding the context window’s capacity does not automatically teach a model how to use it effectively. Therefore, task-driven training curricula are central to developing this capability. These curricula often include tasks like “needle-in-a-haystack” extraction and cross-document reasoning, which explicitly incentivize the model to attend to distant information. Reports on large models such as Qwen2.5-1M (Team, 2025a) and Gemini2.5 (DeepMind, 2025) confirm that curricula aligned with real-world applications markedly improve performance on ultra-long context tasks. Training typically follows a schedule of increasing difficulty, starting with shorter sequences and gradually extending the length, number of distractors, and reasoning complexity.
@@
The quadratic computational cost of the standard self-attention mechanism makes processing very long sequences prohibitively expensive. This has spurred the development of more efficient attention algorithms. Algorithmic methods reduce this cost through sparse connectivity patterns, as seen in models like Longformer (Beltagy et al., 2020) and BigBird (Zaheer et al., 2020), or through linear and kernelized approximations such as the Performer (Choromanski et al., 2020). In parallel, systems-level optimizations focus on improving hardware utilization. A notable example is FlashAttention (Dao et al., 2022), which preserves exact attention while reducing high-bandwidth memory access through kernel fusion. This line of work has been extended by subsequent improvements like Lightning Attention (Zhang et al., 2024a, b). Additional cache-aware and shape-adaptive variants further refine attention allocation to better match task structure (Duanmu et al., 2024; Moonshot AI, 2025). Together, these advances are shifting long-context capabilities from an exotic feature to a practical default, providing a solid substrate for higher-level algorithms and applications.
@@ 5.3.1. Pipeline-based Paradigm
The pipeline-based paradigm for context management is often referred to as context engineering (Liu et al., 2024), where information is manipulated externally through rule-based or modular workflows before being passed to the model.
@@
In this technical path, context re-structuring involves the static pre-arrangement of input. Techniques include chunking and reordering content based on relevance, such as placing critical information at the beginning or end of a prompt to circumvent the “lost-in-the-middle” effect. Structured prompt templates, which assign fixed positions for instructions, examples, and retrieved evidence, are also common, as seen in systems like MemoChat and MemInsight (Fei, 2023; Amazon Science, 2025). For specialized domains, hierarchical organization may be used to manage complex evidence and resolve conflicts, as demonstrated by HiAgent and AWM (Hu, 2024; Wang et al., 2024c).
@@
When evidence is drawn from heterogeneous sources, retrieval and routing becomes critical. Retrieval systems typically combine sparse and dense methods to ensure both keyword and semantic coverage (Karpukhin et al., 2020). A rule-based routing module then selects the most appropriate information based on user intent or knowledge domain. To increase information density, diversity constraints may be applied to reduce redundancy. This approach is exemplified by systems like COPPER, Memary, and COLA, which offer auditable selection policies (Valmeekam et al., 2024; Julio, 2025; Xia, 2025). Production systems like ZEP and Mem0 further integrate joint sparse-dense retrieval for fast and semantically faithful recall (Zep Inc., 2025; Chhikara et al., 2025). While controllable and auditable, the rigidity of these rule-based policies limits their adaptability, motivating a shift toward hybrid, learnable approaches.
@@ 5.3.2. Hybrid Paradigm
In the hybrid paradigm, context re-structuring evolves from simple information arrangement to dynamic knowledge construction (Wang et al., 2024c). A representative technique is the use of learned memory slots, where a lightweight policy network learns to distill and write key information into a structured, finite set of slots rather than passively receiving raw text (Wang and Yan, 2025; Xu et al., 2025a; Yuen et al., 2025). Another approach uses a fine-tuned model to create associations between disparate pieces of knowledge, effectively synthesizing new, structured evidence. Systems such as EMU, Optimus-1, and Nemori illustrate this synthesis-oriented restructuring (Na et al., 2024; Yin et al., 2024; Unknown, 2025), which has been shown to significantly improve reasoning performance (Xu et al., 2025a; Chhikara et al., 2025).
@@
Retrieval and routing in the hybrid paradigm aim to replace fixed rules with learned policies to enhance adaptability. The evolution proceeds mainly along two paths. The first path is modular optimization, which retains the classic pipeline structure but replaces individual rule-based components (e.g., a retriever or re-ranker) with trainable neural modules. This improves performance on specific tasks while maintaining the overall architecture’s interpretability, as exemplified by RAP, SeCom, M+, TWM, RCR-Router, and Learn-to-Memorize (Panasonic Connect, 2024; Qulox, 2025; Chandy, 2025; Wang et al., 2025a; Fang et al., 2025; Zhang et al., 2025a).
The second path is workflow internalization, towards a deeper integration. The LLM is trained to execute the steps of a predefined workflow, learning to decide when to retrieve, rewrite, or evaluate information. Self-RAG, for instance, introduced reflection tokens that allow the model to trigger retrieval and self-critique its generated content based on the results (Asai et al., 2023). Related approaches like TiM, HippoRAG, MemAgent, and Memory-R1 similarly entwine selection and synthesis within the model’s decision-making process, improving adaptability while preserving auditability through explicit step definitions (Wan, 2023; Yao et al., 2024; Yan, 2025; Yan et al., 2025).
@@ 5.3.3. Model-native Paradigm
The highest level of internalization merges the entire memory process into a unified policy space, where the model learns not only how to reason but also how and when to access and manage its memory. Very recently, MemAct (Zhang et al., 2025c) proposes to reframe context management as a learnable, intrinsic capability. In MemAct, the agent actively curates its own working memory by executing explicit editing operations as part of its unified decision-making process, trained via a novel DCPO learning algorithm to balance memory curation with long-term task objectives. This approach contrasts with methods like CARE, which focus on co-adapting retrieval and reasoning without giving the agent explicit control over its memory state (Ma et al., 2025). In the fully native paradigm, the retriever and generator become deeply coupled, often sharing parameters and enabling end-to-end optimization. External knowledge repositories remain valuable as a factual foundation for compliance, but the model itself assumes full autonomy over their use, marking the final transition from an assembly of external modules to a system with internally coherent capabilities.
@@ 5.4. Long-term Memory
Long-term memory refers to persistent knowledge and experience maintained across sessions or tasks. Its contents can be held in two primary types of carriers: external repositories, such as documents, vector stores, or graphs, and the model’s internal parameters, including adapter modules or the weights of the model itself (Packer et al., 2023; Zheng et al., 2024; Park et al., 2023; Edge et al., 2024b; Wu et al., 2022; Wang et al., 2021; Sheng et al., 2023; Du et al., 2025).
Key properties of long-term memory include low write frequency and strong traceability, which support compliance and versioning. The content is typically not editable online during a task; instead, any online optimization focuses on retrieval and utilization strategies. This paradigm provides a factual basis for an agent’s planning and tool use but can face challenges with cross-source consistency and information timeliness. Storing knowledge directly in model parameters can yield lower latency and deeper integration, but it often comes at the cost of reduced interpretability and an increased risk of interference between stored facts (Zhao et al., 2025b). An overview of methods for long-term memory is summarized in Table 2.
@@ 5.4.1. External Repository as Memory Carrier
When using external repositories, the emphasis shifts from how to store information to how to retrieve it precisely and use it reliably (Izacard et al., 2022; Borgeaud et al., 2021). A common practice is to build hybrid indexes offline by combining inverted indices, dense vectors, and metadata. During inference, multi-path retrieval and re-ranking are used to select and compress candidate evidence. This process is often accompanied by deduplication and consistency checks to ensure the injected evidence is both relevant and traceable (Packer et al., 2023; Zheng et al., 2024). While this approach is robust for tasks requiring deep knowledge from large corpora, injecting excessive or poorly selected information can dilute the model’s attention. Consequently, recent methods have moved from single-path semantic retrieval toward more complex pipelines that incorporate learned re-ranking and delegate decisions about when and what to retrieve to learned policies (Edge et al., 2024b).
@@
Because long-term knowledge sources are often lengthy and heterogeneous, simply expanding the context window is not always effective. Layered summarization and structuring are therefore used to manage information density. Documents can be segmented into structured evidence cards with explicit citations and timestamps for verifiability. This structured layer reduces noise and stabilizes the candidate pool for retrieval systems (Park et al., 2023; Packer et al., 2023). The focus has shifted from generating human-readable text summaries to creating machine-usable evidence objects that integrate tightly with retrieval and routing strategies (Edge et al., 2024b; Izacard et al., 2022).
@@
External Repository RETRO (Borgeaud et al., 2021) Text+Vector Static Chunk Retrieval+Cross-attention Industry No 21.12
Atlas (Izacard et al., 2022) Text Static Embedding KNN Industry No 22.08
Generative Agents (Park et al., 2023) Structured Text Rule-based Summarization Scoring Academia No 23.04
Synapse (Zheng et al., 2024) Trajectory Exemplars Offline curation Embedding KNN Academic No 23.06
MemGPT (Packer et al., 2023) Text + Vector Rule-basedSummarization Embedding KNN + Scoring Academia Yes 23.10
GraphRAG (Edge et al., 2024b) Knowledge Graph Rule-based Summarization Graph Retrieval Industry Yes 24.04
LMLM (Zhao et al., 2025b) Text Static Instruction-Trigger Academia No 25.05
Model Parameters K-Adapter (Wang et al., 2021) Adapters /LoRA Bank Offline Training Rule-based+ Router Academia Yes 21.08
Memorizing Transformers (Wu et al., 2022) kNN Cache+KV Online KV writes Embedding KNN Academia Yes 22.03
S-LoRA (Sheng et al., 2023) Adapters /LoRA Bank Offline Training Rule-based + Router Academia No 23.11
MixLoRA (Li et al., 2024e) Adapters /LoRA Bank Joint Fine-Tuning Router (top k) Academia Yes 24.04
ELDER (Li et al., 2024c) Adapters / LoRA Bank Automatic Offline Training Rule-based + Router Academia No 24.08
MoM (Du et al., 2025) Linear State Layers Automatic Offline Traning Router Academia Yes 25.02
@@ 5.4.2. Model Parameters as Memory Carrier
For knowledge that is stable and frequently reused, writing it directly into the model’s parameters can reduce inference latency and dependence on external stores. In practice, this is often achieved through continual pretraining or instruction distillation, where domain-specific knowledge, API schemas, or reasoning patterns are integrated into the training data. Data replay and regularization techniques are used to mitigate catastrophic forgetting. This allows the internalized knowledge to serve as a fallback when external retrieval fails or returns incomplete information (Zhao et al., 2025b; Izacard et al., 2022; Borgeaud et al., 2021).
@@
To balance the high cost of global internalization with the limited capacity of model editing, lightweight modules can be used. This strategy involves freezing the base model and training or attaching small, specialized components that carry specific knowledge. LoRA-style adapters and adapter banks can encode user preferences or task-specific skills in compact modules, while linear state layers can provide additional internal memory that is routed on a per-input basis (Wang et al., 2021; Sheng et al., 2023; Li et al., 2024c, e; Du et al., 2025). This approach is computationally efficient and allows for per-task or per-domain customization.

### v2

@@ 5.1. Overview
To systematically review this evolution, this section divides memory into two primary layers: short-term memory and long-term memory. Our discussion of short-term memory follows its own developmental trajectory. We begin with techniques for long context, which address the fundamental challenge of processing large volumes of information within a single session (Beltagy et al., 2020; Dao et al., 2022; Zaheer et al., 2020). We then examine context management, a more advanced capability focused on the selective organization and utilization of information to mitigate attention dilution and improve reasoning quality (Asai et al., 2023). In contrast, the section on long-term memory reviews methods for persistently storing knowledge across sessions, either in external repository or within model parameters (Wang et al., 2024c; Xu et al., 2025a; Wang et al., 2025b; Zhong et al., 2023). Notably, since long-term memory content is typically not rewritable online, the paradigm shift in this area focuses more on internalizing retrieval and usage strategies rather than the entire storage process.
@@ 5.2.1. Pipeline-based Paradigm
When the dialogue context far exceeds the model’s budget and the necessary knowledge resides in external stores, RAG dynamically constructs the working context for each turn. Compared with traditional RAG pipelines, which typically treat retrieval as a stateless, per-query operation, conversational variants are session-aware: they employ query expansion to capture dialogue intent and multi-path retrieval with reranking to optimize the evidence set across turns. This approach excels at knowledge-intensive, needle-in-a-haystack problems like log analysis and technical question answering. Various frameworks have demonstrated this approach, from those employing MapReduce paradigms (Harrison, 2 10) to more advanced systems that utilize graph-based structures (Edge et al., 2024) or memory-augmented techniques (Qian et al., 2025b; Tang et al., 2024). A key challenge remains that excessive injection of retrieved information can lead to attention dilution. Therefore, practical deployments often treat long windows and session RAG as complementary, using the window to cover retrieval misses and falling back to precise injection when the window becomes too noisy.
@@ 5.2.2. Model-native Paradigm
Expanding the context window’s capacity does not automatically teach a model how to use it effectively. Therefore, task-driven training curricula are central to developing this capability. These curricula often include tasks like “needle-in-a-haystack” extraction and cross-document reasoning, which explicitly incentivize the model to attend to distant information. Reports on large models such as Qwen2.5-1M (Yang et al., 2025e) and Gemini2.5 (Comanici et al., 2025) confirm that curricula aligned with real-world applications markedly improve performance on ultra-long context tasks. Training typically follows a schedule of increasing difficulty, starting with shorter sequences and gradually extending the length, number of distractors, and reasoning complexity.
@@
The quadratic computational cost of the standard self-attention mechanism makes processing very long sequences prohibitively expensive. This has spurred the development of more efficient attention algorithms. Algorithmic methods reduce this cost through sparse connectivity patterns, as seen in models like Longformer (Beltagy et al., 2020) and BigBird (Zaheer et al., 2020), or through linear and kernelized approximations such as the Performer (Choromanski et al., 2022). In parallel, systems-level optimizations focus on improving hardware utilization. A notable example is FlashAttention (Dao et al., 2022), which preserves exact attention while reducing high-bandwidth memory access through kernel fusion. This line of work has been extended by subsequent improvements like Lightning Attention (Qin et al., 2024b, a). Additional cache-aware and shape-adaptive variants further refine attention allocation to better match task structure (Duanmu et al., 2024; Lu et al., 2025d). Together, these advances are shifting long-context capabilities from an exotic feature to a practical default, providing a solid substrate for higher-level algorithms and applications.
@@ 5.3.1. Pipeline-based Paradigm
The pipeline-based paradigm for context management is often referred to as context engineering (Mei et al., 2025c), where information is manipulated externally through rule-based or modular workflows before being passed to the model.
@@
In this technical path, context re-structuring involves the static pre-arrangement of input. Techniques include chunking and reordering content based on relevance, such as placing critical information at the beginning or end of a prompt to circumvent the “lost-in-the-middle” effect. Structured prompt templates, which assign fixed positions for instructions, examples, and retrieved evidence, are also common, as seen in systems like MemoChat and MemInsight (Lu et al., 2023a; Salama et al., 2025). For specialized domains, hierarchical organization may be used to manage complex evidence and resolve conflicts, as demonstrated by HiAgent and AWM (Hu, 2024; Wang et al., 2024c).
@@
When evidence is drawn from heterogeneous sources, retrieval and routing becomes critical. Retrieval systems typically combine sparse and dense methods to ensure both keyword and semantic coverage (Karpukhin et al., 2020). A rule-based routing module then selects the most appropriate information based on user intent or knowledge domain. To increase information density, diversity constraints may be applied to reduce redundancy. This approach is exemplified by systems like Memary, and COLA, which offer auditable selection policies (Julio, 2024; Zhao et al., 2025a). Production systems like ZEP and Mem0 further integrate joint sparse-dense retrieval for fast and semantically faithful recall (Rasmussen et al., 2025; Chhikara et al., 2025). While controllable and auditable, the rigidity of these rule-based policies limits their adaptability, motivating a shift toward hybrid, learnable approaches.
@@ 5.3.2. Hybrid Paradigm
In the hybrid paradigm, context re-structuring evolves from simple information arrangement to dynamic knowledge construction (Wang et al., 2024c). A representative technique is the use of learned memory slots, where a lightweight policy network learns to distill and write key information into a structured, finite set of slots rather than passively receiving raw text (Zhang et al., 2025c; Xu et al., 2025a; Yuen et al., 2025). Another approach uses a fine-tuned model to create associations between disparate pieces of knowledge, effectively synthesizing new, structured evidence. Systems such as EMU, Optimus-1, and Nemori illustrate this synthesis-oriented restructuring (Li et al., 2024e; Unknown, 2025), which has been shown to significantly improve reasoning performance (Xu et al., 2025a; Chhikara et al., 2025).
@@
Retrieval and routing in the hybrid paradigm aim to replace fixed rules with learned policies to enhance adaptability. The evolution proceeds mainly along two paths. The first path is modular optimization, which retains the classic pipeline structure but replaces individual rule-based components (e.g., a retriever or re-ranker) with trainable neural modules. Retrieval-Augmented Planning(RAP) learns to pull and reuse relevant past trajectories instead of following hand-written heuristics (Kagaya et al., 2024), SeCom segments and compresses dialogue into structured memory units and retrieves at the segment level, which reduces noise compared to naive last-k recall in long conversations  (Pan et al., 2025). Temporal Working Memory(TWM) uses query-guided retention over long multimodal streams, keeping only temporally relevant segments rather than forwarding the entire recent history  (Diao et al., 2025). RCR-Router learns routing policies that deliver different memory to different agent or roles instead of broadcasting the same context to everyone  (Liu et al., 2025b). Learn-to-Memorize directly optimizes what to store and what to recall as part of the task objective rather than treating storage and retrieval as fixed utilities (Zhang et al., 2025a). EMU shows the same tendency in multi-agent reinforcement learning by learning which episodic memory are worth recalling to coordinate future behavior  (Na et al., 2024). These works improve performance on specific tasks while maintaining the overall architecture’s interpretability.
The second path is workflow internalization, towards a deeper integration. The LLM is trained to execute the steps of a predefined workflow, learning to decide when to retrieve, rewrite, or evaluate information. Self-RAG, for instance, introduced reflection tokens that allow the model to trigger retrieval and self-critique its generated content based on the results (Asai et al., 2023). HippoRAG integrates retrieval with a structured memory index and incremental consolidation, allowing new information to be linked to existing stored knowledge during generation rather than only before generation  (Gutiérrez et al., 2025). MemAgent treats long context as an active working memory management problem that the model ingests long inputs in stages, refreshes what it keeps, and continues reasoning without relying only on a fixed sliding window (Yu et al., 2025a). Related approaches like TiM and Memory-R1 similarly entwine selection and synthesis within the model’s decision-making process, improving adaptability while preserving auditability through explicit step definitions (tim2023memoryr1_2025).
@@ 5.3.3. Model-native Paradigm
The highest level of internalization merges the entire memory process into a unified policy space, where the model learns not only how to reason but also how and when to access and manage its memory. Very recently, MemAct (Zhang et al., 2025d) proposes to reframe context management as a learnable, intrinsic capability. In MemAct, the agent actively curates its own working memory by executing explicit editing operations as part of its unified decision-making process, trained via a novel DCPO learning algorithm to balance memory curation with long-term task objectives. This approach contrasts with methods like CARE, which focus on co-adapting retrieval and reasoning without giving the agent explicit control over its memory state (Wang et al., 2025f). In the fully native paradigm, the retriever and generator become deeply coupled, often sharing parameters and enabling end-to-end optimization. External knowledge repositories remain valuable as a factual foundation for compliance, but the model itself assumes full autonomy over their use, marking the final transition from an assembly of external modules to a system with internally coherent capabilities.
@@ 5.4. Long-term Memory
Long-term memory refers to persistent knowledge and experience maintained across sessions or tasks. Its contents can be held in two primary types of carriers: external repositories, such as documents, vector stores, or graphs, and the model’s internal parameters, including adapter modules or the weights of the model itself (Packer et al., 2024; Zheng et al., 2024; Park et al., 2023; Edge et al., 2024; Wu et al., 2022; Wang et al., 2021; Sheng et al., 2023; Du et al., 2025).
Key properties of long-term memory include low write frequency and strong traceability, which support compliance and versioning. The content is typically not editable online during a task; instead, any online optimization focuses on retrieval and utilization strategies. This paradigm provides a factual basis for an agent’s planning and tool use but can face challenges with cross-source consistency and information timeliness. Storing knowledge directly in model parameters can yield lower latency and deeper integration, but it often comes at the cost of reduced interpretability and an increased risk of interference between stored facts (Zhao et al., 2025c). An overview of methods for long-term memory is summarized in Table 1.
@@ 5.4.1. External Repository as Memory Carrier
When using external repositories, the emphasis shifts from how to store information to how to retrieve it precisely and use it reliably (Izacard et al., 2022; Borgeaud et al., 2021). A common practice is to build hybrid indexes offline by combining inverted indices, dense vectors, and metadata. During inference, multi-path retrieval and re-ranking are used to select and compress candidate evidence. This process is often accompanied by deduplication and consistency checks to ensure the injected evidence is both relevant and traceable (Packer et al., 2024; Zheng et al., 2024). While this approach is robust for tasks requiring deep knowledge from large corpora, injecting excessive or poorly selected information can dilute the model’s attention. Consequently, recent methods have moved from single-path semantic retrieval toward more complex pipelines that incorporate learned re-ranking and delegate decisions about when and what to retrieve to learned policies (Edge et al., 2024).
@@
Because long-term knowledge sources are often lengthy and heterogeneous, simply expanding the context window is not always effective. Layered summarization and structuring are therefore used to manage information density. Documents can be segmented into structured evidence cards with explicit citations and timestamps for verifiability. This structured layer reduces noise and stabilizes the candidate pool for retrieval systems (Park et al., 2023; Packer et al., 2024). The focus has shifted from generating human-readable text summaries to creating machine-usable evidence objects that integrate tightly with retrieval and routing strategies (Edge et al., 2024; Izacard et al., 2022).
@@
External Repository RETRO (Borgeaud et al., 2021) Text+Vector Static Attention-based Industry No 21.12
Atlas (Izacard et al., 2022) Text Static Embedding-based Industry No 22.08
Generative Agents (Park et al., 2023) Structured Text Rule-based(Summarization) Scoring Academia No 23.04
Synapse (Zheng et al., 2024) Trajectory Exemplars Offline (curation) Embedding-based Academic No 23.06
MemGPT (Packer et al., 2024) Text + Vector Rule-based(Summarization) Embedding + Scoring Academia Yes 23.10
GraphRAG (Edge et al., 2024) Knowledge Graph Rule-based(Summarization) Graph-based Industry Yes 24.04
LMLM (Zhao et al., 2025c) Text Static Instruction-based Academia No 25.05
Model Parameters K-Adapter (Wang et al., 2021) Adapters /LoRA Bank Offline (Training) Rule-based+ Router Academia Yes 21.08
Memorizing Transformers (Wu et al., 2022) kNN Cache+KV Online (KV writes) Embedding-based Academia Yes 22.03
S-LoRA (Sheng et al., 2023) Adapters /LoRA Bank Offline (Training) Rule-based + Router Academia No 23.11
MixLoRA (Li et al., 2024d) Adapters /LoRA Bank Offline (Joint Fine-Tuning) Router (top k) Academia Yes 24.04
ELDER (Li et al., 2025j) Adapters / LoRA Bank Offline (Automatic Training) Rule-based + Router Academia No 24.08
MoM (Du et al., 2025) Linear State Layers Offline (Automatic Training) Router Academia Yes 25.02
@@ 5.4.2. Model Parameters as Memory Carrier
For knowledge that is stable and frequently reused, writing it directly into the model’s parameters can reduce inference latency and dependence on external stores. In practice, Continual pretraining and instruction distillation have been used to inject domain-specific knowledge, API schemas, or reasoning patterns into the base model through training data, as in systems that jointly train a generator and retriever for knowledge-intensive QA (Atlas) or couple parametric knowledge with retrieval supervision in large memory language models (LMLM)  (Zhao et al., 2025c; Izacard et al., 2022). RETRO further shows that some knowledge can remain external but still be treated as internal by conditioning generation on retrieved evidence during training and inference, which lowers the need to scale parameters aggressively  (Borgeaud et al., 2021). Data replay and regularization techniques are used to mitigate catastrophic forgetting. This allows the internalized knowledge to serve as a fallback when external retrieval fails or returns incomplete information.
@@
To balance the high cost of global internalization with the limited capacity of model editing, lightweight modules can be used. This strategy involves freezing the base model and training or attaching small, specialized components that carry specific knowledge. Adapter style methods can encode user preferences or task-specific skills in compact modules, for example K-Adapter for factual andlinguistic knowledge (Wang et al., 2021), or S-LoRA for User preference and skill personalization (Sheng et al., 2023). Recent variants organize many such modules or linear states as banks or mixtures and select them on a per-input basis (Li et al., 2025j, 2024d; Du et al., 2025). These approaches are computationally efficient and allow for per-task or per-domain customization.